A compound contains 59.4% carbon, 10.9% hydrogen, 13.9% nitrogen and 15.8% oxygen, by mass. Find the empirical formula of the compound.


(show working)
1
Expert's answer
2017-10-12T02:52:07-0400
Total percents are:
59.4+10.9+15.8+13.9=100

Atoms ratio can be calculated as
C:H:O:N=59.4/12.0107:10.9/1.0079:15.8/15.994:13.9/14.0067=4.95:10.81:0.988:0.992=5:11:1:1
Thus, empirical formula of the compound is C5H11NO